OpenStudy (anonymous):

For which values of m and n does the graph of f(x)=mx+n intersect the graph of g(x)=1/x in exactly one point and also contain the point (-1,1)?

jimthompson5910 (jim_thompson5910):

f(x)=mx+n has the point (-1,1) on it when f(-1) = 1 f(x)=mx+n f(-1)=m(-1)+n ... replace x with -1 f(-1)=-m+n 1=-m+n ... replace f(-1) with 1 m=n making sense so far?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I wouldn't use point slope form would I?

jimthompson5910 (jim_thompson5910):

not really

jimthompson5910 (jim_thompson5910):

I guess you could, but I don't see how it comes into play

OpenStudy (anonymous):

Wait, why did you replace f(-1) with 1?

jimthompson5910 (jim_thompson5910):

because f(-1) = 1

jimthompson5910 (jim_thompson5910):

the point (-1,1) lies on f(x)

OpenStudy (anonymous):

Okay I get up to the 1= -m+n. But how didi you get rid of the 1?

jimthompson5910 (jim_thompson5910):

oh my bad, typo

jimthompson5910 (jim_thompson5910):

1=-m+n 1+m=n n = m+1

jimthompson5910 (jim_thompson5910):

ready for the next part?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

We're looking for intersection points between f and g. So we set the two functions equal to each other and solve for x. In that process, we'll need to get everything to one side f(x) = g(x) mx+n = 1/x x*(mx+n) = 1 mx^2+nx = 1 mx^2+nx-1 = 0

OpenStudy (anonymous):

Oh seriously. I thought we knew what the y intercept was so we could just plug it into the equation and use trial and error

jimthompson5910 (jim_thompson5910):

no we don't know the y-intercept of f(x)

jimthompson5910 (jim_thompson5910):

we want to force f and g to have exactly one intersection point that means we want to force mx^2+nx-1 = 0 to have exactly one solution

OpenStudy (anonymous):

Okay, so would I factor? but the variables are different though

jimthompson5910 (jim_thompson5910):

The equation mx^2+nx-1 = 0 has exactly one solution when the discriminant D = b^2 - 4ac is equal to zero (ie D = 0) In this case, a = m, b = n, c = -1 D = b^2 - 4ac 0 = b^2 - 4ac ... plug in D = 0 0 = n^2 - 4m*(-1) ... plug in a = m, b = n, c = -1 0 = n^2 + 4m 0 = (m+1)^2 + 4m ... plug in n = m+1

jimthompson5910 (jim_thompson5910):

at this point, you need to solve 0 = (m+1)^2 + 4m for m. You'll get 2 solutions

OpenStudy (anonymous):

would using the quadratic formula work too. I'm not used to the discriminant

jimthompson5910 (jim_thompson5910):

the discriminant is part of the quadratic formula

jimthompson5910 (jim_thompson5910):

the b^2 - 4ac is under the square root

OpenStudy (anonymous):

While solving that for m, I foiled the (m+1)^2 and after adding it to 4m, got m^2+6m+1 I'm not sure that's factorable

jimthompson5910 (jim_thompson5910):

you'll have to use the quadratic formula

jimthompson5910 (jim_thompson5910):

to solve m^2+6m+1 = 0

OpenStudy (anonymous):

Okay, I know how do to it from there. But why did you plug in n=m+1 for that step 0 = n^2 + 4m 0 = (m+1)^2 + 4m ... plug in n = m+1.

OpenStudy (anonymous):

Oh to get like terms, gotcha

jimthompson5910 (jim_thompson5910):

yeah I wanted one variable to solve for